# Single-Channel EEG Tokenization Through Time-Frequency Modeling

### v1

@@
Abstract.
We introduce TFM-Tokenizer, a novel tokenization framework tailored for EEG analysis that transforms continuous, noisy brain signals into a sequence of discrete, well-represented tokens for various EEG tasks. Conventional approaches typically rely on continuous embeddings and inter-channel dependencies, which are limited in capturing inherent EEG features such as temporally unpredictable patterns and diverse oscillatory waveforms. In contrast, we hypothesize that critical time-frequency features can be effectively captured from a single channel. By learning tokens that encapsulate these intrinsic patterns within a single channel, our approach yields a scalable tokenizer adaptable across diverse EEG settings. We integrate the TFM-Tokenizer with a transformer-based TFM-Encoder, leveraging established pretraining techniques from natural language processing—such as masked token prediction—followed by downstream fine-tuning for various EEG tasks. Experiments across four EEG datasets show that TFM-Token outperforms state-of-the-art methods. On TUEV, our approach improves balanced accuracy and Cohen’s Kappa by $5\%$ over baselines. Comprehensive analysis of the learned tokens demonstrates their ability to capture class-distinctive features, enhance frequency representation, and ability to encode time-frequency motifs into distinct tokens, improving interpretability.
EEG, Tokenization, Deep learning
journalyear: 2025copyright: acmcopyrightconference: 2025; ; ccs: Computing methodologies Machine learning approachesccs: Applied computing Health care information systems
1. Introduction
Electroencephalograms (EEGs) reveal real-time neuronal activity with millisecond precision, reflecting the responses to various event stimuli. This makes EEGs essential not only for fundamental research (Conrad et al., 2019; Li et al., 2021) but also for diverse applications such as sleep staging (Phan et al., 2022; Yang et al., 2023; Pradeepkumar et al., 2024), neurological disease detection (Kotoge et al., 2024; Afzal et al., 2024), and brain-computer interfaces (Precup and Bachman, 2012; Wolf et al., 2022; Li et al., 2023). Practical EEG analysis is a complex, multi-step process involving filtering, feature engineering, pattern recognition, annotation, etc. Deep learning (DL) models have shown remarkable success in automating EEG analysis across various tasks (Sharma et al., 2022; Tang et al., 2022; Chen et al., 2023). These achievements stem from their representation learning capability, projecting noisy EEGs onto a discriminative feature space that captures associations with discrete neurophysiological events. Recent neuroscientific literature reports that DL methods successfully match the reliability and accuracy of 20 clinical experts in identifying seizure-related events (Jing et al., 2023).
Figure 1. TFM-Tokenizer tokenizes EEG signals by encoding time-frequency motifs as discrete tokens. The neural tokenizer used for LaBraM (Jiang et al., 2024), which is a raw signal-based tokenizer, fails to capture time-frequency variations. Our TFM-Tokenizer effectively preserves this information.
Despite their success, effectively representing EEGs remains a primary focus in the research community. Real-world EEG recordings are acquired under diverse scenarios, using different devices, and even for the same task, data formats often vary due to mismatched channels and variable lengths (Yang et al., 2024). Unfortunately, most existing methods typically learn representations on a case-by-case basis with specific architectures or fixed channel settings. These methods exhibit limited generalization across tasks and poor scalability to different data formats. For example, a sleep apnea model is not readily applicable to a sleep staging task, since the former aims to capture anomaly patterns while the latter is to learn discriminative representations (Pradeepkumar et al., 2024). Likewise, a sleep staging model trained on central region EEGs cannot directly generalize to frontal EEG data. As a result, retraining, elaborate fine-tuning, or model adjustments are often required (Yang et al., 2024), escalating practical costs and further restricting applicability to advanced devices or non-standard data, such as ECoG (Pradeepkumar et al., 2021). There is thus an urgent need to develop an EEG analysis method that serves broader research objectives.
Recently, the transformative impact of large foundation models(Touvron et al., 2023; Achiam et al., 2023) has inspired and elevated EEG representation learning to new heights. Several foundation EEG models have been proposed (Duan et al., 2023; Yi et al., 2024; Mohammadi Foumani et al., 2024; Yang et al., 2024), demonstrating both enhanced performance and generalization. Researchers often tokenize EEGs into discrete, short-duration snapshots across different data formats and model their dependencies leveraging foundational yet powerful Transformers. By leveraging the paradigm of pretraining on large-scale, cross-task datasets, these models can learn universal representation, enabling effective adaptation to downstream tasks without requiring task-specific architectures (Kotoge et al., 2024; Mohammadi Foumani et al., 2024). Thanks to unified tokenization and the automated input adaptation of the Transformer, such methods also address challenges such as complex data format reconfigurations (Yang et al., 2024) and labeled data scarcity (Jiang et al., 2024). However, this direction remains nascent, and several limitations remain:
  • Inappropriate Tokenization Representation. One reason large language models (LLMs) succeed is their effective tokenization and similar benefits have been demonstrated in image (Van Den Oord et al., 2017) and video (Wang et al., 2024b; Agarwal et al., 2025) tokenization. However, existing foundation EEG models generally do not adopt a discrete tokenization paradigm. Although some methods claim to provide an EEG “tokenizer,” they typically lack a discrete approach similar to NLP. For instance, LaBraM (Jiang et al., 2024) employs a neural tokenizer only during pretraining but relies on raw EEG signals during inference. BIOT (Yang et al., 2024) suggests a “biosignal tokenizer” that remains continuous rather than discretized. Although time-series tokenization methods have shown promise (Ansari et al., 2024; Rasul et al., 2023), they do not scale well to EEG’s higher sampling rates and other artifacts.
  • Insufficient Frequency Representation. Capturing eventful EEG features, which are characterized by distinct frequencies, is a primary focus of EEG analysis. However, tokenizing and modeling raw EEGs often lead to a loss of frequency diversity. This frequency representation collapse is a common issue in time-series modeling, as low-frequency components typically dominate the EEG data, biasing models toward lower frequencies while overlooking critical high-frequency features, such as spikes. Although LaBraM has recently incorporated frequency reconstruction into its objective function, it still relies on raw EEG signals, inevitably resulting in sub-optimal representation, as shown in Figure 1.
  • Scalability and Generalization. EEG-related tasks vary significantly in channel configurations. For example, seizure detection typically utilizes 16 channels, whereas sleep studies often require only 1–2 channels. However, existing models are primarily designed for multi-channel settings, heavily relying on cross-channel prediction. This design limits their scalability and adaptability to configurations with fewer or even single channels, as well as to varying acquisition setups.
Therefore, in this paper, we propose TFM-Token, an effective, fully discretized EEG tokenization framework that captures time-frequency motifs from single-channel EEG signals into distinct tokens. Technically, our contributions are as follows:
  • TFM-Tokenizer and TFM-Encoder: We introduce a scalable discrete tokenization framework for EEG, transforming single-channel signals into discrete token sequences akin to NLP models. TFM-Tokenizer converts EEG into discrete tokens, and TFM-Encoder utilizes them for downstream tasks.
  • Joint Modeling of Frequency and Temporal Dynamics: Our tokenizer integrates raw EEG patches with time-frequency representations, using frequency band and temporal masking to capture essential frequency patterns while disentangling temporal variations (Figure 1).
  • Scable tokenization: Our single-channel approach enables flexible adaptation across EEG tasks and channel configurations. TFM-Tokenizer further enhances state-of-the-art EEG models, such as LaBraM (Jiang et al., 2024).
  • Empirical Validation and Token Quality Analysis: We evaluate our framework on four EEG downstream tasks, demonstrating state-of-the-art performance. Beyond performance, we comprehensively analyze token quality, including token visualization, class-specific uniqueness, and frequency learning analysis, validating that our learned tokens are informative and interpretable.
2. Related work
EEG Representation Learning. To learn general representations and address issues of label scarcity in EEG data, self-supervised learning (SSL) has emerged as a prominent paradigm, and existing works can be categorized into two main approaches: contrastive learning and self-prediction. Contrastive learning methods, including TS-TCC (Eldele et al., 2021), TF-C (Zhang et al., 2022) and SplitSEE (Kotoge et al., 2024), leverage augmentation or transformation of EEG inputs to learn consistent representations. In contrast, self-prediction methods, namely BENDR (Kostas et al., 2021), MAEEG (Chien et al., 2022), BIOT (Yang et al., 2024) and EEG2REP (Mohammadi Foumani et al., 2024), aim to accurately reconstruct masked or corrupted input. However, their learning objectives heavily rely on cross-channel prediction to focus on spatial characteristics. In contrast, our method emphasizes inherent time-frequency features within a single-channel setting and can adapt to any channel configuration.
Foundation EEG Models. Inspired by the success of foundation models in NLP, recent efforts have sought to develop foundation models for EEG analysis. These models can be categorized into decoding and encoder-based methods. Decoding-only methods focus on generative tasks like EEG-to-text translation, with representative works including DeWave (Duan et al., 2023), EEG2Text (Liu et al., 2024), and E2T-PTR (Wang et al., 2024a). In contrast, encoder-only methods concentrate on fundamental EEG classification tasks and representation learning. Notable models include LaBraM (Jiang et al., 2024), BIOT (Yang et al., 2024), BRANT (Zhang et al., 2024), and MMM (Yi et al., 2024). Our work aligns with this latter category, focusing on enhancing the representation quality to improve classification performance.
EEG Tokenization. Tokenization has been instrumental in NLP, where discrete subword units have proven to reduce data complexity and improve model performance and interoperability. Existing attempts for EEGs include patch-based continuous tokenization, such as BIOT (Yang et al., 2024) and BRANT (Zhang et al., 2024), and vector quantization (VQ)-based methods like DeWave (Duan et al., 2023). Patch-based methods do not involve encoding or quantization, leading to unbounded and continuous representations that lack distinctiveness and interpretability. In contrast, VQ-based tokenizers, traditionally successful in tokenizing continuous images (Esser et al., 2020), have recently been adapted for EEG by LaBraM (Jiang et al., 2024), However, its tokenizer is utilized only during training and not during inference. Conceptually, its primary role is to pre-train classification layers, rather than encoding inputs and reducing data complexity. Here, our method is explicitly VQ-based, treating the codebook as a real tokenizer for EEG data. Moreover, we enforce each token to capture time-frequency motifs(Xu et al., 2023) in EEG inputs, ensuring a more structured and interpretable representation.
Frequency Representation Collapse. Frequency domain analysis is crucial in EEG and general time series analysis (Elvander and Jakobsson, 2020; Wu et al., 2021, 2023; Woo et al., 2022). In real-world signals, time-domain observations inherently mix multiple frequency components, and high-energy, low-frequency signals often dominate the spectrum (Huang Norden E Shen Zheng and H, 1998; Lai et al., 2018). As a result, these entangled frequency features makes it difficult for models to distinguish between them (Zhou et al., 2022; Piao et al., 2024). Recent studies have shown that these entangled signals can lead to a collapse in the learned frequency representations (Zhi-Qin John Xu et al., 2020; Piao et al., 2024). Models tend to overemphasize the dominant low-frequency features while neglecting the high-frequency details. This issue can lead to a lack of capturing various EEG waveforms and degenerating data representation (Park and Kim, 2022). Motivated by these works, our paper focuses on developing methods to learn diverse, informative frequency features. In Section 3, we provide an analysis of our proposed frequency-domain tokenizer and its impact on model performance.
3. Preliminary

### v2

# Tokenizing Single-Channel EEG with Time-Frequency Motif Learning

@@
Abstract
Foundation models are reshaping EEG analysis, yet an important problem of EEG tokenization remains a challenge. This paper presents TFM-Tokenizer, a novel tokenization framework that learns a vocabulary of time-frequency motifs from single-channel EEG signals and encodes them into discrete tokens. We propose a dual-path architecture with time–frequency masking to capture robust motif representations, and it is model-agnostic, supporting both lightweight transformers and existing foundation models for downstream tasks. Our study demonstrates three key benefits: Accuracy: Experiments on four diverse EEG benchmarks demonstrate consistent performance gains across both single- and multi-dataset pretraining settings, achieving up to $17\%$ improvement in Cohen’s Kappa over strong baselines. Generalization: Moreover, as a plug-and-play component, it consistently boosts the performance of diverse foundation models, including BIOT and LaBraM. Scalability: By operating at the single-channel level rather than relying on the strict 10–20 EEG system, our method has the potential to be device-agnostic. Experiments on ear-EEG sleep staging, which differs from the pretraining data in signal format, channel configuration, recording device, and task, show that our tokenizer outperforms baselines by $14\%$. A comprehensive token analysis reveals strong class-discriminative, frequency-aware, and consistent structure, enabling improved representation quality and interpretability. Code is available at https://github.com/Jathurshan0330/TFM-Tokenizer.
1 Introduction
Foundation models have revolutionized how machines understand human language, leading to major breakthroughs in natural language processing (NLP) (OpenAI et al., 2024; DeepSeek-AI et al., 2025) and cross-modality tasks such as text-to-image generation (Bordes et al., 2024). Inspired by this success, researchers are now advancing a paradigm shift in electroencephalogram (EEG) analysis toward task-agnostic foundation models (Mohammadi Foumani et al., 2024; Yang et al., 2024; Jiang et al., 2024b; Wang et al., 2024a). By pretraining on massive, diverse EEG data corpora, these models learn universal representations that generalize well across various downstream tasks.
Despite substantial recent progress, an important open problem remains: how to design an effective tokenization method for EEG signals. Tokenization, a core component in NLP, transforms raw text into meaningful tokens, which reduces data complexity and introduces a helpful inductive bias in foundation models (Gastaldi et al., 2025). Typically, tokenization is performed by a learnable function that trains a vocabulary of tokens and statistics from a given corpus. However, existing EEG foundation models tokenize signals by directly segmenting continuous EEGs into short-duration tokens, without learning a vocabulary. They merely discretize EEG signals, failing to capture statistically grounded representations in a data-driven manner. LaBraM (Jiang et al., 2024b) proposes a neural tokenizer to learn data-driven tokens before pretraining. However, these tokens primarily serve as training objectives rather than as actual inputs for subsequent model training and are discarded during downstream inference, limiting their reusability. As a result, the foundation model is still trained on continuous segment-level embeddings, failing to fully leverage the benefits of tokenization, such as improving the quality of input representations. In this paper, we study a novel and critical problem of developing a principled EEG tokenization that seamlessly integrates with various foundation models and enhance downstream performance and generalization.
Figure 1: (a) Our TFM-Tokenizer converts single-channel EEG into discrete tokens by capturing time-frequency motifs. (b) It is adaptable to any different multi-channel settings, (c) can be integrated with existing foundation models to enhance their performance, and (d) enables cross-device scalability.
Various studies have shown that developing an effective tokenization is a non-trivial task in general, as it is influenced by multiple factors (Schmidt et al., 2024). In this paper, we recognize and focus on three key challenges of EEG tokenization. 1) Tokenization target: real-world EEG recordings exhibit diverse formats due to varying devices, channel configurations, and recording lengths (Yang et al., 2024). We argue that tokenizers should be trained and operated at the single-channel level to learn channel-agnostic discrete tokens. This design enables flexible adaptation to multi-channel tasks and can generalize to non-standard EEG devices. In Section 4.4, we provide scalability experiments on ear-EEG settings. 2) Token resolution: in NLP, tokenization can be defined at different resolutions (characters, subwords, words), each reflecting different assumptions about semantic granularity. However, EEG signals are characterized by diverse oscillatory (e.g., alpha, beta) (Pradeepkumar et al., 2024) and transient patterns (e.g., spikes) (Chen et al., 2022). Thus, effective tokens must represent such underlying motifs (Xu et al., 2023) that reflect distinct neural or physiological events. However, these motifs are often distorted by noise, amplitude scaling, and temporal warping, making it challenging to design robust EEG tokenization methods. 3) Tokenization learning objective: EEGs exhibit various temporal variations, manifested as a mixture of low- and high-frequency components that co-occur and are intermixed in complex ways. Relying solely on capturing time‑based motifs into discrete tokens risks losing important spectral structure. We therefore argue that the tokenization learning objective should incorporate time–frequency representations, enabling tokens to encode more meaningful EEG motifs.
To tackle these challenges, we propose TFM-Tokenizer, a novel EEG tokenization framework that captures time–frequency motifs from single-channel EEG signals and encodes them into distinct tokens. Specifically, 1) Tokenizing EEGs at single-channel: We tokenize single-channel EEG signals into discrete token sequences akin to NLP models, which are then paired with a generic transformer to perform multi-channel modeling using these single-channel tokens. Our tokenizer is model-agnostic and can be paired with any downstream model. Our experiments confirmed that TFM-Tokenizer can seamlessly integrate with existing foundation models, and further improve their performance (see Figure 1). 2) Learning motif features as tokens: We introduce a motif learning architecture that encodes time–frequency motifs into tokens through a dual-path encoding design. Capturing frequency-band characteristics or compositions is crucial for EEG analysis, and to model such dynamics, we designed a Localized Spectral Window Encoder, which isolates and aggregates information across frequency bands prior to fusion with temporal features. 3) Explicit time-frequency masking prediction: this learning objective disentangles the entangled time–frequency representations, enabling the model to explicitly learn distinct frequency-specific patterns across time. By forcing the model to predict masked regions in both domains, it encourages the tokenizer to discover and encode meaningful neural motifs that are localized in time and frequency. Overall, our contributions are summarized as follows:
  • Formulating Single-Channel EEG Tokenization. To our knowledge, we are the first to investigate the problem of learning a discrete token vocabulary that captures time–frequency motifs in single-channel EEG signals from a given corpus and directly utilizes them as inputs for downstream modeling.
  • Proposing Novel TFM-Token Framework. We introduce a single-channel EEG tokenization framework that transforms EEG into a discrete token sequence via TFM-Tokenizer, which is then used by a lightweight transformer model for cross-channel and downstream modeling. As shown in Figure 1c, TFM-Tokenizer integrates smoothly with existing models and consistently boosts performance, improving BIOT and LaBraM by approximately $4\%$ on TUEV dataset.
  • Broad Evaluation across Foundation Models and Devices. Extensive experiments across four datasets show that our method outperforms strong baselines, achieving up to a $17\%$ gain over the baseline model on TUEV dataset. We also evaluate cross-device scalability on an ear-EEG sleep staging task, using electrodes outside the standard 10–20 EEG system, where our tokenizer outperforms baselines by $14\%$. Beyond performance, we comprehensively analyze token quality, including token consistency, class-specific uniqueness, and frequency learning analysis, validating that our learned tokens are informative and interpretable.
2 Related Work
EEG Foundation Models and Tokenization Methods. Existing EEG foundation models can be categorized into decoding and encoder-based methods. Decoding-based methods focus on generative tasks like cross-modal translation  (Duan et al., 2023; Liu et al., 2024; Wang et al., 2024c). In contrast, encoder-based methods focus on classification tasks and representation learning. Notable models include LaBraM  (Jiang et al., 2024b), BIOT  (Yang et al., 2024), BRANT  (Zhang et al., 2024), and MMM  (Yi et al., 2024). Our work aligns with this latter category, aiming to enhance input representations to improve classification performance and generalization across diverse foundation models. A parallel question is how to tokenize EEG signals. Existing methods primarily adopt segment-based continuous tokenization (Yang et al., 2024; Wang et al., 2024b; Zhang et al., 2024). Vector Quantized (VQ) tokenizers (Van Den Oord et al., 2017), which have been successful in tokenizing continuous images (Esser et al., 2020), have recently been adapted for EEG by LaBraM (Jiang et al., 2024b). However, in LaBraM, the tokenizer is not designed to represent EEG data and replace raw signals as inputs to foundation models; instead, it mainly serves as a training objective. In this paper, we propose a new tokenization framework for EEG signals that encodes inputs into discrete representations and provide a reusable interface for foundation models.
EEG Motif Learning. Motifs are short, recurring patterns with small variability in a time series and may hold predictive or discriminative value (Xu et al., 2023). In the EEG domain, motif learning remains largely underexplored, with only a few works such as (Schäfer & Leser, 2022), which focus solely on the temporal domain. EEG motifs correspond to neurophysiological events such as oscillatory bursts or transient spikes, which are best characterized by joint temporal-spectral structure. Frequency-domain modeling is therefore essential, yet raw time-domain signals often entangle multiple spectral components. This can cause models to overemphasize dominant low-frequency rhythms while overlooking informative high-frequency details (Zhi-Qin John Xu et al., 2020; Piao et al., 2024). Such bias limits the ability to capture diverse EEG waveforms and degrades representation quality (Park & Kim, 2022). To the best of our knowledge, we are the first to propose methods to encode diverse, informative time–frequency motifs as discrete tokens.
3 Methodology
3.1 Framework Overview and Forward Process
Our TFM-Tokenizer framework consists of two major phase, as shown in Figure 2:
  1. 1.
    TFM-Tokenizer with Motif Learning. The tokenizer is trained in a single-channel, unsupervised setting, capturing key motif features. We regard motifs as various waveforms that encode characteristic time–frequency patterns in EEGs. To represent these motifs, the tokenizer is composed of four components: (i) a Localized Spectral Window Encoder that extracts frequency patterns within short spectral windows, (ii) a Temporal Encoder that incorporates raw EEG context, (iii) a Temporal Transformer that models dependencies across windows, and (iv) a codebook quantizer that maps embeddings into a discrete vocabulary. Therefore, we train a motif-based vocabulary that transforms continuous EEGs into interpretable discrete tokens (Sec. 3.2).
  2. 2.
    Downstream Transformer Model. This phase serves as an example to illustrate how a foundation model processes tokenized sequences for downstream tasks such as classification. Raw EEGs are first passed through our pretrained tokenizer, where they are converted into discrete tokens that serve as inputs to foundation models. Since the tokenizer is model-agnostic, it can be paired with different backbone models. In our implementation, we adopt a lightweight Transformer (Vaswani, 2017) with linear attention (Katharopoulos et al., 2020), demonstrating that the tokenizer ($\sim$0.7M parameters) enables strong performance even with a compact model (Sec. 3.3).
Overall, we first pretrain the tokenizer to learn a discrete vocabulary of EEG motifs. The tokenizer is then frozen, and the downstream Transformer is pretrained with a masked token prediction objective. Finally, the downstream Transformer is fine-tuned on target EEG tasks such as classification.